# Energy-Efficient Designs for SIM-Based Broadcast MIMO Systems

### v1

@@ Abstract
Stacked intelligent metasurface (SIM), which consists of multiple layers of intelligent metasurfaces, is emerging as a promising solution for future wireless communication systems. In this timely context, we focus on broadcast multiple-input multiple-output (MIMO) systems and aim to characterize their energy efficiency (EE) performance. To gain a comprehensive understanding of the potential of SIM, we consider both dirty paper coding (DPC) and linear precoding and formulate the corresponding EE maximization problems. For DPC, we employ the broadcast channel (BC)- multiple-access channel (MAC) duality to obtain an equivalent problem, and optimize users’ covariance matrices using the successive convex approximation (SCA) method, which is based on a tight lower bound of the achievable sum-rate, in combination with Dinkelbach’s method. Since optimizing the phase shifts of the SIM meta-elements is an optimization problem of extremely large size, we adopt a conventional projected gradient-based method for its simplicity. A similar approach is derived for the case of linear precoding. Simulation results show that the proposed optimization methods for the considered SIM-based systems can significantly improve the EE, compared to the conventional counterparts. Also, we demonstrate that the number of SIM meta-elements and their distribution across the SIM layers have a significant impact on both the achievable sum-rate and EE performance.
Index Terms:
Optimization, broadcast, EE, MIMO, stacked intelligent metasurface (SIM), multi-user.
@@ I Introduction
The framework for the future development of International Mobile Telecommunications (IMT) for 2030 highlights sustainability as a fundamental goal for future communication systems [1]. This means that these systems are expected to be designed with minimal environmental impact, focusing on the efficient use of resources, reducing power consumption, and lowering greenhouse gas emissions. Due to this, the study and development of energy-efficient wireless communications have recently attracted much of attention. At the same time, the global mobile network data traffic is expected to reach 563 exabytes (EBs) by 2029 [2]. To accommodate such a high volume of data traffic, existing network technologies need to evolve, providing additional capabilities. For example, conventional MIMO systems are advancing toward massive MIMO (mMIMO) and ultra-massive MIMO (umMIMO) systems. However, a large number of radio frequency (RF) chains required to support mMIMO transmissions results in substantial power consumption, which leads to an unsustainable and energy inefficient communication model.
@@
A promising technical solution that addresses the growing demand for higher data rates while simultaneously enhancing EE is based on the use of intelligent metasurfaces, specifically reconfigurable intelligent surfaces. RISs are composed of a large number of programmable metamaterial or tiny discrete antenna elements. Each of these elements is capable, using integrated electronic circuits, of dynamically adjusting its electromagnetic (EM) properties (i.e., to form EM fields with controllable amplitudes, phases, polarization) and consequently its EM response to the incoming waves. In this way, RISs can modify the incoming waves in a programmable and controllable manner [3]. This capability allows RISs to simultaneously improve multiple performance metrics, such as spectrum efficiency, EE, coverage. Unfortunately, the multiplicative effect of the path loss of the RIS-assisted links significantly limits the potential EE gains from RIS deployment.
@@
To address the critical issue of RIS, several innovative approaches have been proposed that utilize metamaterial-based antenna technologies instead of conventional antenna arrays in mMIMO transceiver design. These include holographic radio, dynamic metasurface antennas and SIMs. Holographic radio, also known as holographic MIMO (HMIMO), is a hybrid transceiver architecture that achieves high directive gain, spectral efficiency and EE by incorporating a continuous structure of densely packing sub-wavelength metamaterial antenna elements. These element, combined with holographic techniques, are capable of recording and reconstructing the amplitude and the phase of wave fronts [4]. The significantly lower power consumption of HMIMO allows for the deployment of a greater number of antenna elements compared to traditional mMIMO, resulting in higher EE [5]. Similarly, DMAs consist of multiple microstrips, each composed of a multitude of metamaterial radiating elements and connected to a single RF chain [6]. Due to this, DMAs achieve better EE performance than even hybrid analog-to-digital (A/D) architectures, since they do not need additional power to support numerous phase shifters [7]. However, both HMIMOs and DMAs are single layer matasuface structures, which may require a very large number of elements due to practical hardware constraints that limit the number of tunable amplitudes/phases associated with each meta-element.
In contrast, SIMs represent the latest advancement in metamaterial-based antenna technologies. SIMs consist of multiple parallel metasurface layers, each accommodating numerous meta-elements with programmable phase characteristics. These layers are integrated with conventional radio transceivers that employ a small to moderate number of active antennas. The concept of SIMs draws inspiration from the architecture of a deep neural network (DNN), which is a multi-layer neuron structure capable of implementing various functions [8]. Similarly, SIMs can efficiently implement different signal processing tasks, such as transmit precoding and receive combining, directly in the EM domain when properly controlled and programmed. Hence, SIMs have the potential to substantially improve the performance metrics of conventional communication systems, such as the achievable rate and the EE, while requiring minimal additional hardware complexity.
In [9], SIMs were exploited to implement a 2D discrete Fourier transform (DFT) for direction of arrival direction of arrival (DOA) estimation. Moreover, a hybrid channel estimator was proposed in [10], in which the received training symbols were initially processed in the wave domain and subsequently in the digital domain. In [11], the authors jointly optimized the transmit beamforming at the base station (BS) and the SIM phase shifts, to minimize the Cramer-Rao bound (CRB) for target estimation. Using an experimental SIM platform, they evaluated the performance of the proposed algorithms for communication and sensing tasks.
A general path loss model for an SIM-assisted wireless communication system was developed in [12], based on which, an algorithm aimed at maximizing the received power was derived. In [13], the authors studied the achievable sum-rate maximization problem for a downlink channel between a SIM-assisted BS and multiple single-antenna users. The achievable rate optimization for a downlink multi-user SIM-assisted system using statistical channel state information (CSI) was proposed in [14]. Utilizing statistical CSI, the ergodic sum-rate was optimized for a satellite communication system in [15]. In [16], a joint optimization of the SIM phase shifts and transmit power allocation for maximizing the sum-rate in a SIM-assisted multi-user multiple-input single-output (MISO) communication system was implemented, employing a deep reinforcement learning (DRL) approach. In [17], the authors optimized the achievable rate in an uplink SIM-based cell-free MIMO architecture with distributed signal processing. In this setup, each access point (AP) performs local detection of user information, and a central processing unit (CPU) subsequently combines these local estimates to recover the final user information.
@@
The integration of SIMs with transmitters and receivers into a so-called SIM-based HMIMO system, which performs signal precoding and combining in the wave domain, was proposed in [18]. The introduced channel fitting approach enables the SIM-based HMIMO system to achieve significant channel capacity gains compared to mMIMO and RIS-assisted counterparts. Furthermore, the optimization of achievable rates for the SIM-based HMIMO system was studied in [19]. An approach for the mutual information maximization in a SIM-based HMIMO system with discrete signaling was presented in [20], using the cutoff rate as an alternative metric. This study demonstrates that incorporating even a small-scale digital precoder into the system can substantially increase the mutual information performance.
@@
Despite the extensive research summarized in the aforementioned papers, the EE analysis of SIM-assisted MIMO systems remains unexplored. Motivated by this gap, we aim to maximize the EE for a SIM-aided broadcast system with DPC and linear precoding. To find the maximum achievable EE in the case of DPC, we formulate a joint optimization problem of the covariance matrix of the transmitted signal and the phase shifts of the SIM meta-elements. For linear precoding, which is more practical, we consider a joint optimization problem of the transmit signal precoding and the phase shifts of the SIM meta-elements. In both cases, the BS has a limited total power budget and the SIM meta-elements are subject to the unit modulus constraint. The main contributions of this paper are summarized as follows:
@@
  • For DPC, we exploit the well-known Gaussian MIMO BC-MAC duality, reformulating the original EEmax problem as a function of the users’ covariance matrices in the MAC and the phase shifts of the SIM meta-elements. In the context of the adopted AO framework, we present an efficient solution for optimizing the users’ covariance matrices. This solution is based on a tight and concave lower bound of the achievable sum-rate, which is derived using the SCA method. By applying Dinkelbach’s method, we then obtain the optimal users’ covariance matrices by closed-form expressions. Our complexity analysis demonstrates that our proposed method has significantly lower complexity compared to an existing solution. For the optimization of the phase shifts of the SIM meta-elements, we employ a conventional projected gradient-based method, updating all SIM layers in parallel. This approach is viable considering the large size of this problem. In this context, we derive closed-form expressions for the complex-valued gradients involved.
@@
  • For linear precoding, we leverage an interesting recent result for the sum-rate maximization that allows for reformulating the considered EEmax problem as an equivalent one, but with a greatly reduced dimension. After this important step, we again invoke the SCA method to derive a quadratic lower bound of the achievable sum-rate and approximate the EEmax problem as a concave fractional program. Next, we apply Dinkebach’s method to solve the resulting problem, where optimal users’ precoders are found by closed-form expressions. Similar to the DPC-based scheme, the phase shifts of the SIM meta-elements in this setting are optimized in parallel using a conventional projected gradient-based method.
@@
  • We present efficient implementations of the proposed algorithms, analyze their computational complexities in terms of the number of complex multiplications, and mathematically prove their convergence.
@@
  • We show through simulation results that the proposed algorithms can substantially increase the EE in SIM-aided broadcast communication systems, with greater improvements observed in the case of DPC. Moreover, we demonstrate that using the aforementioned precoding schemes is crucial to mitigate the impact of multi-user inference, especially in systems with a large number of users. We also provide several valuable insights into the design and performance of SIM-based holographic MIMO systems. First, we show that the EE is highly dependent on the number and distribution of SIM meta-elements across the SIM layers. Second, we find that the EE for a SIM-aided system with a low number of meta-elements can even be lower than the EE for a conventional MIMO system without SIM integration. Third, in SIM-aided broadcast systems without digital precoding, optimal EE transmission involves activating only a subset of the available transmit antennas, where each antenna in this subset transmits an independent data stream. Lastly, we demonstrate that at least 3 bits per meta-element are required to ensure that the reduction in EE caused by quantization errors remains within acceptable limits.
Notation: Bold lower and upper case letters represent vectors and matrices, respectively. $\mathbb{C}^{m\times n}$ denotes the space of $m\times n$ complex matrices. $\mathbf{H}^{T}$ and $\mathbf{H}^{H}$ denote the transpose and Hermitian transpose of $\mathbf{H}$, respectively. $|\mathbf{H}|$ is the determinant of $\mathbf{H}$ and $\operatorname{Tr}(\mathbf{H})$ denotes the trace of $\mathbf{H}$. $\log_{2}(\cdot)$ is the binary logarithm, $\ln(\cdot)$ is the natural logarithm, $(\cdot)^{+}$ denotes the pseudo-inverse and $\left(\cdot\right)^{\ast}$ denotes the complex conjugate. $\left\|\mathbf{H}\right\|$ denotes the Frobenius norm of $\mathbf{H}$ which reduces to the Euclidean norm if $\mathbf{H}$ is a vector. $\operatorname{vec}_{d}(\mathbf{H})$ is the vector comprised of the diagonal elements of $\mathbf{H}$. The notation $\mathbf{A}\succeq(\succ)\mathbf{B}$ means that $\mathbf{A}-\mathbf{B}$ is positive semidefinite (definite). $\mathbf{I}$ represents an identity matrix whose size depends from the context. $\Re(\mathbf{x})$ and $\Im(\mathbf{x})$ denote the real and imaginary part of $\mathbf{x}$, respectively. For a vector $\mathbf{x}$, $\operatorname{diag}(\mathbf{x})$ denotes a diagonal matrix with the elements of $\mathbf{x}$ on the diagonal. $\mathcal{CN}(\mu,\sigma^{2}$) denotes a circularly symmetric complex Gaussian random variable with mean $\mu$ and variance $\sigma^{2}$. $|x|$ denotes the modulus of the complex number $x$, and |$\mathbf{H}$| denotes the determinant of $\mathbf{H}$. Finally, we denote by $\nabla_{\mathbf{x}}f(\cdot)$ the complex gradient of $f(\cdot)$ with respect to (w.r.t.) $\mathbf{x}^{\ast}$, i.e., $\nabla_{\mathbf{x}}f(\cdot)=\frac{1}{2}\Bigl{(}\frac{\partial f(\cdot)}{\partial\Re(\mathbf{x})}+j\frac{\partial f(\cdot)}{\partial\Im(\mathbf{x})}\Bigr{)}$.
@@ II-A System Model
We consider a multi-user broadcast system in which a BS with $N_{t}$ transmit antennas communicates with $K$ users, where each user has $N_{r}$ receive antennas. The BS is also equipped with a SIM, which consists of $L$ metasurface layers with $N$ meta-elements per layer. In general, SIMs are controlled by external field programmable gate array (FPGA) devices, which adjust the phase shifts of individual meta-elements, thereby implementing signal beamforming directly in the EM wave domain. [^1] By considering both SIM (i.e., wave-based precoding) and digital precoding, our system model is general enough to include the wave-based only precoding as a special case.
The phase shifts of the meta-elements in the $l$-th SIM layer are presented by the diagonal matrix $\boldsymbol{\mathbf{\Phi}}^{l}=\text{diag}(\boldsymbol{\phi}^{l})=\text{diag}([\phi_{1}^{l}\>\phi_{2}^{l}\>\cdots\>\phi_{N}^{l}]^{T})$, where $\phi_{n}^{l}=\exp(j\theta_{n}^{l})$ and $\theta_{n}^{l}$ is the phase shift introduced by the $n$-th element of the $l$-th layer. Signal propagation between two consecutive layers, $l-1$ and $l$, of the SIM is modeled by the matrix $\mathbf{W}^{l}\in\mathbb{C}^{N\times N}$ for $l=2,3,\dots,L$. More precisely, signal propagation between the $n$-th meta-element of the $(l-1)$-th and the $m$-th meta-element of $l$-th layer of the SIM is presented by the $(m,n)$-th element of $\mathbf{W}^{l}$, which is calculated according to the Rayleigh-Sommerfeld diffraction theory as [21, Eq. (1)]
$[\mathbf{W}^{l}]_{m,n}=\frac{A_{k}\cos\chi_{m,n}}{d_{m,n}}\Bigl{(}\frac{1}{2\pi d_{m,n}}-\frac{j}{\lambda}\Bigr{)}e^{j\frac{2\pi d_{m,n}}{\lambda}}$ (1)
where $A_{k}$ is the area of each meta-element, $d_{m,n}$ is the distance between the meta-elements of these two layers of the SIM, $\chi_{m,n}$ is the angle between the propagation direction and normal direction of the $(l-1)$-th layer, and $\lambda$ is the wavelength. Signal propagation between the transmit antenna array and the first layer of the SIM is presented by the matrix $\mathbf{W}^{1}\in\mathbb{C}^{N\times N_{t}}$, whose elements can be calculated as in (1). Finally, the EM response of the transmit SIM can be written as
$\mathbf{B}=\boldsymbol{\mathbf{\Phi}}^{L}\mathbf{W}^{L}\cdots\boldsymbol{\mathbf{\Phi}}^{2}\mathbf{W}^{2}\boldsymbol{\mathbf{\Phi}}^{1}\mathbf{W}^{1}\in\mathbb{C}^{N\times N_{t}}.$ (2)
For the considered system, the end-to-end channel matrix between the BS and the $k$-th user receive antenna array is given by
@@
where $\mathbf{G}_{k}$ $\in\mathbb{C}^{N_{r}\times N}$ denotes the channel matrix between the final layer of the SIM and user $k$. We assume that $\mathbf{G}_{k}$ is perfectly known to the BS in an effort to investigate its full theoretical potential.

### v2

@@ Abstract
Stacked intelligent metasurface (SIM), which consists of multiple layers of intelligent metasurfaces, is emerging as a promising solution for future wireless communication systems. In this timely context, we focus on broadcast multiple-input multiple-output (MIMO) systems and aim to characterize their energy efficiency (EE) performance. To explore the potential of SIM, we consider both dirty paper coding (DPC) and linear precoding (LP) and formulate the corresponding EE maximization problems. For DPC, we employ the broadcast channel (BC)- multiple-access channel (MAC) duality to obtain an equivalent problem, and optimize users’ covariance matrices using the successive convex approximation (SCA) and Dinkelbach’s methods. Since the phase shift optimization problem of the SIM meta-elements is one of extremely large size, we adopt a conventional projected gradient-based method for its simplicity. A similar approach is followed for the case of LP. Simulation results show that the proposed optimization methods for the considered SIM-based systems can significantly improve the EE, compared to conventional counterparts. Also, we demonstrate that the number of SIM meta-elements and their distribution across the SIM layers have a significant impact on both the achievable sum-rate and EE performance.
@@ I Introduction
The framework for the future development of International Mobile Telecommunications (IMT) for 2030 highlights sustainability as a fundamental goal for future communication systems, with the aim of minimizing environmental impact through efficient resource usage, reduced power consumption, and lower greenhouse gas emissions [1]. This has made energy-efficient wireless communications a key research focus. At the same time, global mobile network data traffic is projected to reach 563 exabytes (EBs) by 2029 [2], which requires the evolution of network technologies to meet the increasing demand. For example, conventional MIMO systems are advancing toward massive MIMO (mMIMO) and ultra-massive MIMO (umMIMO) systems. However, these technologies face challenges in EE due to the substantial power consumption required to support a large number of radio frequency (RF) chains.
@@
A promising solution to address these challenges is the use of reconfigurable intelligent surfaces, which consist of a large number of programmable metamaterial elements capable of dynamically tuning their electromagnetic (EM) properties. In this way, RISs can modify the incoming waves in a programmable and controllable manner [3]. This capability allows RISs to simultaneously improve multiple performance metrics, such as spectrum efficiency, EE, and coverage. Unfortunately, the significant path-loss of RIS-assisted links greatly limits the potential EE gains from RIS deployment.
@@
To address the limitations of RIS, several innovative metamaterial-based antenna technologies such as holographic radio, dynamic metasurface antennas, and SIMs have been proposed. Holographic radio, also known as holographic MIMO (HMIMO), achieves high directive gain, spectral efficiency (SE) and EE by combining densely packing sub-wavelength metamaterial antenna elements with holographic techniques[4, 5]. Similarly, DMAs consist of multiple microstrips, each composed of a multitude of metamaterial radiating elements and connected to a single RF chain [6]. This design offers better EE than hybrid analog-to-digital (A/D) architectures by eliminating the need for additional power to support numerous phase shifters [7]. However, both HMIMOs and DMAs, as single-layer structures, often require a very large number of elements due to hardware constraints that limit the number of tunable amplitudes/phases.
In contrast, SIMs represent the latest advancement in metamaterial-based antenna technologies. Composed of multiple parallel metasurface layers with nearly passive, programmable meta-elements, SIMs can be integrated with conventional transceivers using a small number of active antennas, offering a low-cost and energy-efficient solution. Inspired by deep neural networks (DNN), SIMs perform signal processing tasks, such as transmit precoding and receive combining, directly in the EM domain when properly controlled and programmed [8]. The close spacing between active antennas and SIM layers significantly reduces the path-loss and eliminates the multiplicative path-loss effects seen in RIS-aided systems. By utilizing whole metasurfaces to achieve array gain and suppress co-channel interference in the EM wave domain, SIMs reduce transmission power requirements while significantly enhancing EE and overall system performance, all with minimal hardware complexity.
Recent studies have investigated various applications of SIMs. In [9], SIMs were exploited to implement a 2D discrete Fourier transform (DFT) for direction of arrival (DOA) estimation. A hybrid channel estimator combining wave and digital domain processing was proposed in [10]. However, it requires perfect orthogonality between user pilot sequences, which is hard to achieve. In [11], the authors minimized the Cramer-Rao bound (CRB) for target estimation in an ISAC system with a SIM. Specifically, using an experimental SIM platform, they showed power variations across layers but without any physical explanation. A general path-loss model for SIM-assisted systems was developed in [12], based on which, algorithms aimed at maximizing the received power were derived. This design can generate complex radiation patterns but requires many layers for discrete phase shifts. Research on sum-rate maximization includes studies on SIM-assisted downlink MISO channels [13] and multi-user systems using statistical channel state information (CSI) [14]. In particular, it was shown that using statistical CSI yields slightly worse performance while achieving lower computational and processing overhead. In [15], deep reinforcement learning (DRL) was shown to outperform conventional optimization approaches for maximizing the sum rate in SIM-assisted communication systems. In uplink SIM-based cell-free MIMO systems, rate optimization using distributed signal processing was studied in [16]. In this scenario, each access point (AP) locally detects user information, which is then combined by the central processing unit (CPU) to recover the final data, though this process can demand substantial link resources in practice.
@@
The integration of SIMs with transmitters and receivers into a SIM-based HMIMO system, capable of performing signal precoding and combining in the wave domain, was proposed in [17]. This system introduced a channel fitting approach that enables the SIM-based HMIMO system to achieve significant channel capacity gains compared to mMIMO and RIS-assisted counterparts. An approach for the mutual information maximization in a SIM-based HMIMO system with discrete signaling was presented in [18], using the cutoff rate as an alternative metric. This study demonstrates that incorporating even a small-scale digital precoder into the system can substantially increase its mutual information.
@@
Despite extensive research on SIM-assisted systems, the EE analysis of SIM-assisted MIMO systems remains largely unexplored. Notably, EE optimization inherently includes SE maximization and transmit power minimization as special cases. Motivated by this gap, we aim to maximize the EE in SIM-aided broadcast systems using DPC and LP. To solve the formulated EE maximization (EEmax) problems, we blend multiple optimization techniques to derive efficient problem-driven algorithms, leveraging the special structure of the considered problems and suitable optimization tools. For DPC, we formulate a joint optimization problem of the covariance matrix of the transmitted signal and the phase shifts of the SIM meta-elements. In the more practical case of LP, we consider a similar joint optimization problem for transmit precoding and SIM phase shifts. For both scenarios, we assume a limited total power budget and impose unit modulus constraints on the SIM meta-elements. Furthermore, perfect CSI is assumed at the BS to explore the full EE potential of SIM-based systems. Note that we do not optimize the number of active transmit antennas, which is left for future work. The main contributions of this paper are summarized as follows:
@@
  • For DPC, we exploit the well-known Gaussian MIMO BC-MAC duality to reformulate the EEmax problem in terms of users’ covariance matrices in the MAC and the phase shifts of the SIM meta-elements. Within the adopted alternating optimization (AO) framework, we present an efficient solution for optimizing the covariance matrices, using a tight concave lower bound of the achievable sum-rate via the SCA method. Dinkelbach’s method is then applied to obtain closed-form expressions for the optimal covariance matrices. Our complexity analysis demonstrates that our proposed method has significantly lower complexity compared to an existing solution [19]. For optimizing the SIM phase shifts, we employ a projected gradient-based method to update all SIM layers in parallel. This approach is viable considering the large size of the considered problem. In this context, we derive closed-form expressions for the complex-valued gradients involved.
@@
  • For LP, we leverage a recent result for the sum-rate maximization that allows for reformulating the considered EEmax problem as an equivalent one, with a greatly reduced dimension. Following this important step, we again invoke the SCA method to derive a quadratic lower bound of the achievable sum-rate and approximate the EEmax problem as a concave fractional program. Next, we apply Dinkelbach’s method to solve the resulting problem, obtaining optimal users’ precoders in closed-form expressions. Similar to the DPC-based scheme, the phase shifts of the SIM meta-elements in this setting are optimized in parallel using a conventional projected gradient-based method.
@@
  • We present efficient implementations of the proposed algorithms, analyze their computational complexity in terms of the number of complex multiplications, and mathematically prove their convergence.
@@
  • We show through simulation results that the proposed algorithms substantially increase the EE in SIM-aided broadcast communication systems, with greater improvements observed for DPC. The results highlight the importance of the proposed precoding schemes in mitigating multi-user interference, especially in systems with a large number of users. We also provide several valuable insights into the design and performance of energy-efficient SIM-based broadcast MIMO systems. First, the EE increases logarithmically with the number of meta-elements per SIM layer, with a nearly double gain when increasing from 49 to 100 compared to 100 to 196. Second, SIM-aided systems with a low number of meta-elements may have lower EE than conventional MIMO systems without SIM integration. Third, a few SIM layers suffice to balance SE gains and power consumption for optimal EE. Fourth, in SIM-aided broadcast systems without digital precoding, optimal EE is achieved by activating a subset of transmit antennas, each transmitting independent data streams. Lastly, at least 3 bits per meta-element are required to keep EE degradation from quantization errors within acceptable limits.
The mathematical notation used in this paper is summarized in Table I.
TABLE I: Notation description.
Symbol Description
$\mathbf{H}$ Matrix
$\mathbf{h}$ Row/column vector
$\mathbb{C}^{m\times n}$ Space of $m\times n$ complex matrices
$\mathbf{H}^{T}$ Transpose of $\mathbf{H}$
$\mathbf{H}^{H}$ Hermitian transpose of $\mathbf{H}$
$|\mathbf{H}|$ Determinant of $\mathbf{H}$
$\operatorname{Tr}(\mathbf{H})$ Trace of $\mathbf{H}$
$\log_{2}(\cdot)$ Binary logarithm
$\ln(\cdot)$ Natural logarithm
$(\cdot)^{+}$ Pseudo-inverse operation
$\left\lfloor\cdot\right\rfloor$ Floor operation
$\left(\cdot\right)^{\ast}$ Complex conjugate
$\left\|\mathbf{H}\right\|$ Frobenius norm
$\operatorname{vec}_{d}(\mathbf{H})$ Vector consisting of the diagonal elements of $\mathbf{H}$
$\mathbf{A}\succeq(\succ)\mathbf{B}$ $\mathbf{A}-\mathbf{B}$ is positive semidefinite (definite)
$\mathbf{I}$ Identity matrix
$\Re(\mathbf{x})$ Real part of $\mathbf{x}$
$\Im(\mathbf{x})$ Imaginary part of $\mathbf{x}$
$\operatorname{diag}(\mathbf{x})$ Matrix with the elements of $\mathbf{x}$ on the main diagonal
$\mathcal{CN}(\cdot,\cdot$) Circularly symmetric complex Gaussian distribution
$|x|$ Modulus of the number $x$
[…]
$\nabla_{\mathbf{x}}f(\cdot)$ Complex gradient of $f(\cdot)$ with respect to (w.r.t.) $\mathbf{x}^{\ast}$
@@ II-A System Model
System level diagrams of conventional and SIM-aided broadcast systems are shown in Fig. 1. In both cases, we consider a multi-user broadcast system in which a BS with $N_{t}$ transmit antennas communicates with $K$ users, each equipped $N_{r}$ receive antennas. The main difference is in the presence of a SIM at the BS, which comprises of $L$ metasurface layers, each containing $N$ meta-elements. The SIM is controlled by external field programmable gate array (FPGA) devices, which adjust the phase shifts of individual meta-elements for wave domain beamforming. [^1] By considering both SIM (i.e., wave-based precoding) and digital precoding, our system model is general enough to include the wave-based only precoding and digital only precoding as special cases. For the sake of simplicity, we assume that this device is integrated into the BS hardware and communicates with the SIM via a dedicated interface.
The phase shifts of the meta-elements in the $l$-th SIM layer are represented by the diagonal matrix $\boldsymbol{\mathbf{\Phi}}^{l}=\text{diag}(\boldsymbol{\phi}^{l})=\text{diag}([\phi_{1}^{l}\>\phi_{2}^{l}\>\cdots\>\phi_{N}^{l}]^{T})$, where $\phi_{n}^{l}=\exp(j\theta_{n}^{l})$ and $\theta_{n}^{l}$ is the phase shift of the $n$-th element in the $l$-th layer. Signal propagation between two consecutive layers, l1l-1 and $l$, is modeled by the matrix $\mathbf{W}^{l}\in\mathbb{C}^{N\times N}$ for $l=2,3,\dots,L$. Specifically, the signal propagation between the $n$-th meta-element of the (l1)(l-1)-th and the $m$-th meta-element of $l$-th layer of the SIM is presented by[20, Eq. (1)]
$\mathbf{W}^{l}]_{m,n}=\frac{A_{k}\cos\chi_{m,n}}{d_{m,n}}\Bigl(\frac{1}{2\pi d_{m,n}}-\frac{j}{\lambda}\Bigr)e^{j\frac{2\pi d_{m,n}}{\lambda}}$ (1)
where $A_{k}$ is the area of each meta-element, $d_{m,n}$ is the distance between the meta-elements of these two layers, $\chi_{m,n}$ is the angle between the propagation direction and the normal to the (l1)(l-1)-th layer, and $\lambda$ is the wavelength. The signal propagation between the transmit antenna array and the first layer of the SIM is represented by the matrix $\mathbf{W}^{1}\in\mathbb{C}^{N\times N_{t}}$, whose elements can be calculated as in (1).
(a) Conventional broadcast system.
(b) SIM-aided broadcast system.
Figure 1: System level diagrams of conventional and SIM-aided broadcast systems.
The end-to-end channel matrix between the BS and the $k$-th user is expressed as
@@
where $\mathbf{G}_{k}$ $\in\mathbb{C}^{N_{r}\times N}$ denotes the channel matrix between the final layer of the SIM and user $k$, and $\mathbf{B}=\boldsymbol{\mathbf{\Phi}}^{L}\mathbf{W}^{L}\cdots\boldsymbol{\mathbf{\Phi}}^{2}\mathbf{W}^{2}\boldsymbol{\mathbf{\Phi}}^{1}\mathbf{W}^{1}\in\mathbb{C}^{N\times N_{t}}$ is the EM response of the transmit SIM. We assume that $\mathbf{G}_{k}$ is perfectly known to the BS to investigate its full theoretical potential.